OpenStudy (anonymous):

PLEASE HELP!!! find the vertex, axis of symmetry, the intercepts, the domain, the range, intervals where the function increases, intervals where the function decreases and graph y=x^2+4x

OpenStudy (anonymous):

where would you like to start?

OpenStudy (anonymous):

the vertex please

OpenStudy (anonymous):

ok the first coordinate of the vertex of \(y=ax^2+bx+c\) is \(-\frac{b}{2a}\) in your case \(a=1,b=4\) so what is the first coordinate of the vertex?

OpenStudy (anonymous):

or -2

OpenStudy (anonymous):

not quite right the second time

OpenStudy (anonymous):

first coordinate of the vertex is \(-2\) and the second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by \(-2\)

OpenStudy (anonymous):

btw finding the vertex answer almost every other question above

OpenStudy (anonymous):

I got a(-2)^2+4(-2)+0

OpenStudy (anonymous):

-2a-8+0 ???

OpenStudy (anonymous):

there is no \(a\) in this , it is just \((-2)^2+4(-2)\)

OpenStudy (anonymous):

which give you \(4-8=-4\) for the \(y\) coordinate

OpenStudy (anonymous):

ok no a because a=1?

OpenStudy (anonymous):

yeah i just wrote \(y=ax^2+bx+c\) as a general form your specific one is \(y=x^2+4x\)

OpenStudy (anonymous):

ok my vertex is (-2,-4)

OpenStudy (anonymous):

right now we can breeze through most of the other questions

OpenStudy (anonymous):

axis of symmetry since the first coordinate of the vertex is \(-2\) the axis of symmetry is \(x=-2\)

OpenStudy (anonymous):

it's that simple?

OpenStudy (anonymous):

the domain since this is a polynomial, the domain is all real numbers

OpenStudy (anonymous):

is the range y

OpenStudy (anonymous):

yeah that simple the range since the second coordinate of the vertex is \(-4\) and this parabola opens up the range is \([-4,\infty)\) or \(y\geq -4\)

OpenStudy (anonymous):

ok did we skip the intercepts?

OpenStudy (anonymous):

intervals where the function decreases since the axis of symmetry is \(x=-2\) it decreases on \((-\infty, -2)\) or \(x<-2\)

OpenStudy (anonymous):

so evidently increases on \(x>-2\) vertex tells all and it is easy skipped intercepts because that is not from the vertex, we have to do something else

OpenStudy (anonymous):

that is why i skipped the intercepts but they are not hard, it is just that they do not come from the vertex

OpenStudy (anonymous):

ok so that's why I got it wrong...please explain

OpenStudy (anonymous):

\(y\) intercept is easiest that comes by putting \(x=0\) and you get if \(x=0\) then \(y=0\) and so the \(y\) intercept is 0

OpenStudy (anonymous):

\(x\) intercepts put \(y=0\) and get \[x^2+4x=0\] then solve for \(x\) which is not hard in this example

OpenStudy (anonymous):

factor as \(x(x+4)=0\)and you see that the zeros are \(x=0\) or \(x=-4\)

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

o I got it...so both 0nd -4 are the answers

OpenStudy (anonymous):

lets go slow we are finding the \(x\) intercepts, where the graph crosses the \(x\) axis

OpenStudy (anonymous):

it crosses the \(x\) axis where \(y=0\) so we set \(x^2+4x=0\) and get \(x(x+4)=0\) making \(x=-4\) or \(x=0\) the \(x\) intercepts, as ordered pairs, are \((-4,0)\) and \((0,0)\)

OpenStudy (anonymous):

and the \(y\) intercept is also \((0,0)\)

OpenStudy (anonymous):

now we got it all, right?

OpenStudy (anonymous):

yes, so my graph will have 3 points

OpenStudy (anonymous):

it is a parabola that opens up the minimum point on the parabola is the vertex, which is at \((-2,-4)\) and it is symmetric about \(x=-2\) and crosses the \(x\) axis at \((-4,0)\) and \((0,0)\) aka the origin

OpenStudy (anonymous):

this is the part that confuses me

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

here is a nice picture

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=x^2%2B4x

OpenStudy (anonymous):

you can see all the feature, decreasing, increasing, vertex, x intercepts etc

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

thanks for the photo this really helps

OpenStudy (anonymous):

good, wolfram is nice to use